# ChatDBG: An AI-Powered Debugging Assistant

### v1

@@ IV-A Using ChatDBG: Preliminaries
Instructions:
  You are a debugging assistant. You will be given a Python stack trace for an error and answer questions related to the root cause of the error. Call the debug function to run Pdb debugger commands on the stopped program. You may call the debug function to run the following commands: bt, up, down, p expression, list. Call debug to print any variable value or expression that you believe may contribute to the error. Call the info function to get the documentation and source code for any variable, function, package, class, method reference, field reference, or dotted reference visible in the current frame. Examples include: n, e.n where e is an expression, and t.n where t is a type. Unless it is from a common, widely-used library, you MUST call info exactly once on any symbol that is referenced in code leading up to the error. Call the provided functions as many times as you would like. The root cause of any error is likely due to a problem in the source code from the user. Explain why each variable contributing to the error has been set to the value that it has. Continue with your explanations until you reach the root cause of the error. Your answer may be as long as necessary. End your answer with a section titled “Recommendation” that contains one of: a fix if you have identified the root cause a numbered list of 1-3 suggestions for how to continue debugging if you have not
Enriched Stack Trace:
  The program has this stack trace: [... skipping 4 hidden frame(s)] ./bootstrap.py(19)<module>()      15      16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag() ---> 19 resampled_stats(observed_marbles, 5)    Global variables:      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’) > ./bootstrap.py(16)resampled_stats()      14                                 num_trials)      15 ---> 16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag()    Variables in this frame:      num_trials: int = 5      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’)      stats: ndarray = array([’B’, R’, B’, ..., R’,        B’, R’], dtype=’<U32’)
Error:
  The program encountered the following error: AssertionError The code assert len(stats) == num_trials is correct and must not be changed.
History:
  This is the history of some pdb commands I ran and the results:   (ChatDBG) p num_trials   5
User Text:
  Why doesn’t stats have 5 elements?
Figure 5: The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
IV-B Debugging a Target Program
@@
IV-C ChatDBG Architecture Overview
ChatDBG orchestrates communication between the user, the debugger, as shown in the architecture diagram and ChatDBG’s top-level command processing loop in Figure 4. The operations in the pseudocode map naturally onto debugger APIs and onto LLM APIs supporting completion and function calls [4]. ChatDBG currently utilizes OpenAI’s API [17] and GPT-4 models. We elaborate on the most salient technical innovations after this overview.
@@ IV-C ChatDBG Architecture Overview
Command Debugger Output
code loc LLDB The source code surrounding loc, where loc is filename:lineno.
definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc is filename:lineno.
TABLE II: ChatDBG command extensions. These commands are available to not only the user but also the LLM, and they provide access to information beyond what is typical in a debugger.
IV-D Initial Prompts and Enriched Stack Traces
In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error. We highlight the components comprising that context in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2.
Instructions. The instructions at the top of the prompt ask the LLM to answer questions about the root cause of the error, to focus on user code, to explain values stored in variables, and to end each response with either a fix or suggestions for further debugging steps. The last item ensures a relatively consistent structure on answers that facilitates reading them and evaluating their quality. Paragraphs 2-4 of the instructions are the take the wheel prompt described in Section IV-E.
@@ IV-D Initial Prompts and Enriched Stack Traces
In Python, ChatDBG leverages Pdb’s internal data structures to build enriched stack traces. When converting values to suitable string representations, ChatDBG must balance utility with the size of the string produced. For objects, ChatDBG calls the object’s __repr__ method if an appropriate (non-default) version exists. Otherwise, it iterates over the object’s fields and recursively converts their values to strings. Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number. The rest of the elements are just abbreviated with an ellipsis (…). This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses.
@@
s11 181 crash Value Error Takes a sample from a table smaller than the sample size without replacement
@@
s14 65 crash Index Error Simulation fails to properly create a random sample under null hypothesis
@@
TABLE III: Python programs exhibiting a variety of common errors. Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code. Semantic errors reflect failed tests expressed as assertions. Crashes reflect unexpected termination due to any other type of error.
IV-E Taking the Wheel
ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models [4]. This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response. The LLM calls these functions by sending special messages to the client as part of its response stream. The client receives those messages, computes the requested results, and sends them back to the LLM. The initial prompt describes how to use the available functions.
IV-F Navigating the Code
While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions. ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code. These commands are available to the LLM via function calls and are listed in Figure II.
The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol. Instead, ChatDBG adds two other debugging commands to LLDB. The first, code, prints the code surrounding a source location described by a filename and line number, as in code polymorph.c:118. The second command, definition, prints the location and source code for the definition corresponding to the first occurrence of a symbol on a given line of code. For example, definition polymorph.c:118 target prints the location and source for the declaration of target corresponding to its use on that line. The definition implementation leverages the clangd language server, which supports source code queries via JSON-RPC and Microsoft’s Language Server Protocol [18].
IV-G Slices for Interactive Python
ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks. Interactive sessions lead to many individual code cells that are each evaluated separately. Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables. Others have noted the challenges of reasoning about program behavior in this context [19, 20]. ChatDBG provides an additional slice debugging command to facilitate that reasoning. The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells. It returns the code for cells in that slice. Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
In[2]: def make_marble_sample(): ...
In[3]: def proportion_blue(sample): ...
In[4]: def resampled_stats(observed_marbles, num_trials):
           stats = bootstrap_statistic(observed_marbles,
                                       proportion_blue,
                                       num_trials)
           assert len(stats) == num_trials
           return stats
In[5]: observed_marbles = make_marble_sample()
       stats = resampled_stats(observed_marbles, 5)
After evaluating these cells, slice(observed_samples) returns the source for the cells labeled In[2] and In[5], and slice(stats) returns the source for all four cells. ChatDBG uses ipyflow to compute slices [20, 21].
V Evaluation
@@
We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs. The former have well-defined expected behavior that enable us to thoroughly and systematically assess ChatDBG. The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes.
Our evaluation addresses the following research questions: RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python? RQ2: Which components of ChatDBG contribute to its effectiveness? RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
V-A Python
@@
We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table III. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks [22] from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
@@
Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions. We ran each program ten times under the five configurations in Table IV: Default Stack includes standard stack traces, as generated by ipdb [23], with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel. Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands. These three configurations all use why? as the initial user text. +Targeted Question asks a question specific to the failure. For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they are intended to represent. For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
c3 (Crash)
@@
ChatDBG used the gpt-4-1106-preview model for these experiments. Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model [24], and a completion time of about 25 seconds. Subsequent steps in extended debugging dialogs incurred comparable costs. Time was highly variable and dominated by the performance of OpenAI’s service. These characteristics will be different for other platforms and models.
RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
Each response was manually examined and deemed a success if it included an accurate explanation of the error and an actionable fix. That fix could be either code or a prose description in which all necessary details were made explicit. Figure 6 shows the success rate under each configuration.
RQ1 Summary: Even with just the simple question why?, ChatDBG was successful 57% of the time. With questions specialized to the target’s particular error, that number jumps to 67%, and with an additional dialog step ChatDBG succeeded in identiyfing and fixing the defect in 85% of the trials.
@@
TABLE IV: Configurations used in the Python experiments.
Figure 6: Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
Figure 7: Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
Program LoC Error Type Root Cause   Proximate Cause Fix Root Cause Fix
BC [25] 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
GZIP [25] 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
NCOM [25] 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
PEG [26] 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
POLY [25] 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
TIFF [26] 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
YAML1 [26] 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
YAML2 [26] 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
TABLE V: Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
@@
Figure 7 presents the success rates for each program under each configuration. The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particuarly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure. However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
@@
The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior. When the LLM is asked to continue its reasoning in +Dialog, ChatDBG’s success rate improves despite the follow-up prompt providing no feedback on the contents or quality of the first response. This phenom indicates that constraints on the underlying LLM’s response lengths may prevent it from conducting the amount of reasoning necessary to successfully develop a fix in a single step. The success rates for +Targeted Question and +Dialog demonstrate the importance of continued dialogs and user input. We expect those features to be even more important to ChatDBG’s success when diagnosing bugs in more complex programs.
The LLM also demonstrated its background knowledge with the responses including, for example, details of Python idioms and libraries, the definition of h-index [27], and the implementation and limitations of various statistical techniques.
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow ups.

### v2

@@ 4.1. Using ChatDBG: Preliminaries
4.2. Debugging a Target Program
@@
[…]
Figure 4. ChatDBG architecture and command processing algorithm.
4.3. ChatDBG Architecture Overview
ChatDBG orchestrates communication between the user, the debugger, and the LLM, as shown in the architecture diagram in Figure 4. The operations in the command loop pseudocode map naturally onto debugger APIs and onto LLM APIs supporting completion and function calls (OpenAI, 2023a). ChatDBG currently utilizes OpenAI’s API (OpenAI, 2020) and GPT-4 models. We provide a brief overview of the ChatDBG architecture and then elaborate on the most salient technical innovations below.
@@ 4.3. ChatDBG Architecture Overview
{justify}
Instructions:
  • You are a debugging assistant. You will be given a Python stack trace for an error and answer questions related to the root cause of the error.
  • Call the debug function to run Pdb debugger commands on the stopped program. You may call the debug function to run the following commands: bt, up, down, p expression, list. Call debug to print any variable value or expression that you believe may contribute to the error.
  • Call the info function to get the documentation and source code for any variable, function, package, class, method reference, field reference, or dotted reference visible in the current frame. Examples include: n, e.n where e is an expression, and t.n where t is a type. Unless it is from a common, widely-used library, you MUST call info exactly once on any symbol that is referenced in code leading up to the error.
  • Call the provided functions as many times as you would like.
  • The root cause of any error is likely due to a problem in the source code from the user. Explain why each variable contributing to the error has been set to the value that it has. Continue with your explanations until you reach the root cause of the error. Your answer may be as long as necessary.
  • End your answer with a section titled “Recommendation” that contains one of:
    • a fix if you have identified the root cause
    • a numbered list of 1-3 suggestions for how to continue debugging if you have not
Enriched Stack Trace:
The program has this stack trace:
[... skipping 4 hidden frame(s)]./bootstrap.py(19)<module>()     15     16     assert len(stats) == num_trials     17     18 observed_marbles = make_marble_bag()---> 19 resampled_stats(observed_marbles, 5)   Global variables:     observed_marbles: ndarray = array([’R’, ’R’, ’R’, ..., ’B’,                                        ’B’, ’B’], dtype=’<U1’)> ./bootstrap.py(16)resampled_stats()     14                                 num_trials)     15---> 16     assert len(stats) == num_trials     17     18 observed_marbles = make_marble_bag()   Variables in this frame:     num_trials: int = 5     observed_marbles: ndarray = array([’R’, ’R’, ’R’, ...,                                        ’B’, ’B’, ’B’], dtype=’<U1’)     stats: ndarray = array([’B’, ’R’, ’B’, ..., ’R’, ’B’,                             ’R’], dtype=’<U32’)
Error:
The program encountered the following error:
AssertionErrorThe code assert len(stats) == num_trials is correct and must not be changed.
History:
This is the history of some pdb commands I ran and the results:
(ChatDBG) p num_trials5
User Text:
Why doesn’t stats have 5 elements?
Figure 5. The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
4.4. Initial Prompts and Enriched Stack Traces
In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error. We illustrate the components of the prompt in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2 as a running example.
Instructions. The instructions at the top of the prompt ask the LLM to answer questions about the root cause of the error, to focus on user code, to explain values stored in variables, and to end each response with either a fix or suggestions for further debugging steps. The last item ensures a relatively consistent structure for answers that facilitates reading them and evaluating their quality. Paragraphs 2-4 of the instructions are the take the wheel prompt described in Section 4.5.
@@ 4.4. Initial Prompts and Enriched Stack Traces
In Python, ChatDBG leverages Pdb’s internal data structures to build enriched stack traces. When converting values to suitable string representations, ChatDBG must balance utility with the size of the string produced. For objects, ChatDBG calls the object’s __repr__ method if an appropriate (non-default) version exists. Otherwise, it iterates over the object’s fields and recursively converts their values to strings. Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number. The rest of the elements are just abbreviated with an ellipsis (…). This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses. This strategy balances the need to provide the LLM with sufficient information to diagnose the error with the need to avoid overwhelming it with too much information. In cases with the elided details are important, the LLM can request them via the take the wheel mechanism.
@@
4.5. Taking the Wheel
ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models (OpenAI, 2023a). This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response. The LLM calls these functions by sending special messages to the client as part of its response stream. The client receives those messages, computes the requested results, and sends them back to the LLM. The initial prompt describes how to use the available functions.
Supporting agentic reasoning over run-time program state via function calls is a key technical innovation of ChatDBG. Without this capability, there would be no effective way to provide the LLM with a detailed view of relevant program state. A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (DBLP:conf/nips/LewisPPPKGKLYR020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval. However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
Command Debugger Output
code loc LLDB The source code surrounding loc, where loc has the form filename:lineno.
definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc has the form filename:lineno.
Table 2. ChatDBG command extensions. ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code. ChatDBG provides access to them via the LLM’s function call API.
4.6. Navigating the Code
While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions. Trying to include all possibly-relevant source code in the initial prompt would be infeasible and would prevent ChatDBG from scaling to larger codebases. Instead, ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code. These commands are available to the LLM via function calls and are listed in Table 2.
The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol. Instead, ChatDBG adds two other debugging commands to LLDB. The first, code, prints the code surrounding a source location described by a filename and line number, as in code polymorph.c:118. The second command, definition, prints the location and source code for the definition corresponding to the first occurrence of a symbol on a given line of code. For example, definition polymorph.c:118 target prints the location and source for the declaration of target corresponding to its use on that line. The definition implementation leverages the clangd language server, which supports source code queries via JSON-RPC and Microsoft’s Language Server Protocol (Microsoft, 2016).
4.7. Slices for Interactive Python
ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks. Interactive sessions lead to many individual code cells that are each evaluated separately. Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables. Others have noted the challenges of reasoning about program behavior in this context (Head et al., 2019; Shankar et al., 2022). ChatDBG provides an additional slice debugging command to facilitate that reasoning. The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells. It returns the code for cells in that slice. Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
[…]
[fontsize=]pythonslice.py
After evaluating these cells, slice(observed_samples) returns the source for the cells labeled In[2] and In[5], and slice(stats) returns the source for all four cells. ChatDBG uses ipyflow to compute slices (Shankar et al., 2022; IPyflow, 2020).
4.8. Security and Risks
It is possible for the LLM to issue debugging commands containing arbitrary code through the debug function call provided by ChatDBG. That code could, for example, delete files or execute other malicious actions on the client. ChatDBG mitigates this risk by sanitizing LLM-generated debugging commands before running them. For Python, the sanitizer ensures that any functions called in LLM-provided commands belong to a user-configurable whitelist. For native code, code provenance is harder to track and languages are more permissive, so the sanitizer rejects any commands calling functions. ChatDBG supports an --unsafe flag to disable sanitizing when the client system is running in an isolated environment that obviates the need for such protections.
It is also possible for the LLM to hallucinate and respond with incorrect or misleading diagnoses and fixes. ChatDBG mitigates this risk by not directly applying proposed code fixes or suggestions to the target code. Instead, ChatDBG presents them to the user, who may then vet and judge the quality of the LLM’s responses and decide whether or not to follow suggested changes.
@@ 4.8. Security and Risks
s11 181 crash Value Error Creates a sample without replacement that is larger than the input
@@
s14 65 crash Index Error Incorrect computation of a random sample under null hypothesis
@@
Table 3. Python programs exhibiting a variety of common errors. Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code. Semantic errors reflect failed tests expressed as assertions. Crashes reflect unexpected termination due to any other type of error.
5. Evaluation
@@
We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs. The former have well-defined expected behavior that enable us to thoroughly and systematically assess ChatDBG. The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes. Our evaluation addresses the following research questions: RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python? RQ2: Which components of ChatDBG contribute to its effectiveness? RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
5.1. Python
@@
We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table 3. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
@@
Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions. We ran each program ten times under the five configurations in Table 4: Default Stack includes standard stack traces, as generated by ipdb (ipdb, 2007), with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel. Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands. These three configurations all use why? as the initial user text. +Targeted Question asks a question specific to the failure. For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they are intended to represent. For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
c3 (Crash):
@@
ChatDBG used the gpt-4-1106-preview model for these experiments. Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, [n. d.]), and a completion time of about 25 seconds. Subsequent steps in extended debugging dialogs incurred comparable costs. Time was highly variable and dominated by the performance of OpenAI’s service. These characteristics will be different for other platforms and models and, given current trends, we expect significant reductions in both time and cost as models improve.
@@
Table 4. Configurations used in the Python experiments.
Figure 6. Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
Each response was manually examined and deemed a success if it included an accurate explanation of the error and an actionable fix. That fix could be either code or a prose description in which all necessary details were made explicit. To avoid bias in this assessment, explicit criteria for each program was determined prior to examining the responses.
Figure 6 shows the success rate under each configuration. With all features enabled, ChatDBG was successful at identifying and fixing bugs in well over half of the trials. Any time or energy expended by the user manually debugging those cases would be all but eliminated by using ChatDBG.
RQ1 Summary: Even with just the simple question why?, ChatDBG was successful 57% of the time. With questions specialized to the target’s particular error, that number jumps to 67%, and with an additional dialog step ChatDBG succeeded in identifying and fixing the defect in 85% of the trials.
@@
Figure 7 presents the success rates for each program under each configuration. The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particularly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure. However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
@@
Figure 7. Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
Error Fix
Program LoC Type Root Cause Proximate Cause Root Cause
BC (Lu et al., 2005) 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
GZIP (Lu et al., 2005) 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
NCOM (Lu et al., 2005) 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
PEG (An et al., 2023) 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
POLY (Lu et al., 2005) 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
TIFF (An et al., 2023) 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
YAML1 (An et al., 2023) 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
YAML2 (An et al., 2023) 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior in queries. When the LLM is asked to continue its reasoning in +Dialog, ChatDBG’s success rate improves despite the follow-up prompt providing no feedback on the contents or quality of the first response. This phenom indicates that constraints on the underlying LLM’s response lengths may prevent it from conducting the amount of reasoning necessary to develop a fix in a single step. The success rates for +Targeted Question and +Dialog demonstrate the importance of continued dialogs and user input. We expect those features to be even more important to ChatDBG’s success when diagnosing bugs in more complex programs.
The LLM also demonstrated its background knowledge with the responses including, for example, details of Python idioms and libraries, the definition of h-index (Hirsch, 2005), and the implementation and limitations of various statistical techniques.
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people!) to find it. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow ups.